OpenStudy (anonymous):

in a rectangle that is 15 x 8 meters, what is the diagonal length from two corners

OpenStudy (kropot72):

Use the theorem of Pythagoras: \[\large Diagonal\ length=\sqrt{15^{2}+8^{2}}\ meters\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

isn't the Pythagorean theorem a^2 + b^2 = c^2 or is square root(a^2 + a^2) just a simpler version of doing that

OpenStudy (anonymous):

*a^2 + b^2

OpenStudy (kropot72):

We used the theorem to find the square of the diagonal length. To finish we need to take the square root.